OpenStudy (anonymous):

What is the written form of the fraction 1/5 ? Samuel added 1/5 gallon of gas to the lawn mower. A. one fifth B. one-fifth

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

Incorrect.

OpenStudy (anonymous):

1/5 = one fifth 1/25 = one twenty-fifth 2/100,000 = two hundred-thousandths 200/1000 = two hundred thousandths

OpenStudy (anonymous):

Oh! I get it! Thanks!

OpenStudy (anonymous):

My pleasure. ^_^